OpenStudy (anonymous):

approximately how much air does it take to fill up the beach ball below if it has an 18-inch radius

OpenStudy (anonymous):



OpenStudy (ddcamp):

V = (4/3)(pi)(r³) Though "how much air" is a somewhat ambiguous question, I'm assuming they just want cubic inches.

OpenStudy (anonymous):

yes they want cubic inches

OpenStudy (anonymous):

You're looking for how much air the ball can hold, so you'd be looking for the volume. The formula for the volume of a sphere is: \[\frac{ 4 }{ 3 }\pi r^3\] So just plug into the equation, r being 18.

OpenStudy (anonymous):

i got 24416.64

OpenStudy (anonymous):

That's the same answer I got. So your answer would be 24416.64in^3

OpenStudy (anonymous):

but the answers they got is 5832 cubic inches, 18,322 cubic inches, 24,429 cubic inches, 73,287 cubic inches

OpenStudy (anonymous):

its either one of those

OpenStudy (anonymous):

It depends on how you plugged it in, AI guess. I just tried putting it into my calculator again as (4/3)(\[(4/3)(\pi)(18^3)\]

OpenStudy (anonymous):

I ended up with 24429 Cubic Inches